a bus traveled on leveled road for 2 hours at an average speed of 20 miles per hour faster than it did on the winding road. the time sent on the winding road was 5 hours. find the average speed of the level road if the entire trip was 320 miles
Answer by ptfile(81)

Let r+20 = the speed of the bus on leveled road
r = the speed of the bus on winding road
2hrs = the duration of the bus traveled on leveled road
5hrs = the duration of the bus traveled on winding road
A bus traveled on leveled road for 2 hours at an average speed of 20 miles per hour faster than it did on the winding road. The time sent on the winding road was 5 hours. Find the average speed of the level road if the entire trip was 320 miles.
Distance = (rate)(time)
2(r+20)+ 5r = 320
2r+40+5r = 320
7r = 280
r = 40
Substitute r with 40
40 + 20 = 60
The bus's speed on the level road is 60mph.